# Enhancing Ethereum Fraud Detection via Generative and Contrastive Self-supervision

### v1

@@ Abstract
The rampant fraudulent activities on Ethereum hinder the healthy development of the blockchain ecosystem, necessitating the reinforcement of regulations. However, multiple imbalances involving account interaction frequencies and interaction types in the Ethereum transaction environment pose significant challenges to data mining-based fraud detection research. To address this, we first propose the concept of meta-interactions to refine interaction behaviors in Ethereum, and based on this, we present a dual self-supervision enhanced Ethereum fraud detection framework, named Meta-IFD. This framework initially introduces a generative self-supervision mechanism to augment the interaction features of accounts, followed by a contrastive self-supervision mechanism to differentiate various behavior patterns, and ultimately characterizes the behavioral representations of accounts and mines potential fraud risks through multi-view interaction feature learning. Extensive experiments on real Ethereum datasets demonstrate the effectiveness and superiority of our framework in detecting common Ethereum fraud behaviors such as Ponzi schemes and phishing scams. Additionally, the generative module can effectively alleviate the interaction distribution imbalance in Ethereum data, while the contrastive module significantly enhances the framework’s ability to distinguish different behavior patterns. The source code will be released on GitHub soon.
@@ I Introducing
Figure 1: Illustration of the imbalanced distribution in account interaction frequency.
@@ II-B Graph-based Fraud Detection
Furthermore, the homogeneous transaction graphs constructed by the aforementioned methods neglect the heterogeneity of accounts and transaction behaviors. To address this, Zhou et al. [5] incorporate call records from contract accounts as features for externally owned accounts and propose a hierarchical attention encoder based on contrastive learning to effectively learn account representations, alleviating the issue of label scarcity. Du et al. [34] utilize a similar transaction graph construction method and design a GNN-based link prediction method to de-anonymize accounts in mixing services. Jin et al. [35, 36] predefine static and temporal metapaths to describe the complex interaction behaviors in Ethereum and design a metapath-based account feature augmentation module, which effectively improves the performance of existing Ponzi scheme detection methods. Considering that defining metapaths requires expert knowledge and is time-consuming, Liu et al. [37] and Xu et al. [38] employ transformer models to autonomously acquire metapath patterns, subsequently aggregating neighbor accounts based on these metapaths to obtain higher-order information for enhancing fraud detection.
@@ III-B Graph Modeling for Ethereum Fraud Detection
Figure 3: Illustration of the imbalanced distribution in account interaction types.
@@ III-C Refinement of Ethereum Interactions
Unlike the coarse-grained interactions trans and call which only reflect the ways in which accounts interact with each other, meta-interactions further describe the ways in which specific types of accounts can interact with each other, and thus can be referred to as fine-grained interactions. Based on the definition of meta-interaction and the characteristics of Ethereum, we can identify six types of meta-interactions from HEIG, as shown in Fig. 2(b):
$r^{1}_{\text{ca}}:\text{CA}\stackrel{{\scriptstyle\text{{call}}}}{{\longrightarrow}}\text{CA},\quad\ \ r^{1}_{\text{eoa}}:\text{EOA}\stackrel{{\scriptstyle\text{{call}}}}{{\longrightarrow}}\text{CA}$ (1)
$r^{2}_{\text{ca}}:\text{CA}\stackrel{{\scriptstyle\text{{trans}}}}{{\longrightarrow}}\text{CA},\quad\ \ r^{2}_{\text{eoa}}:\text{EOA}\stackrel{{\scriptstyle\text{{trans}}}}{{\longrightarrow}}\text{CA}$
$r^{3}_{\text{ca}}:\text{CA}\stackrel{{\scriptstyle\text{{trans}}}}{{\longrightarrow}}\text{EOA},\quad r^{3}_{\text{eoa}}:\text{EOA}\stackrel{{\scriptstyle\text{{trans}}}}{{\longrightarrow}}\text{EOA}$
Furthermore, we define =$\mathcal{R}=\{r^{1}_{\text{ca}},r^{2}_{\text{ca}},r^{3}_{\text{ca}},r^{1}_{\text{eoa}},r^{2}_{\text{eoa}},r^{3}_{\text{eoa}}\}$ to denote the set of meta-interactions.
@@ IV-A Fine-grained Ethereum Interaction Feature Generation
In complex Ethereum interaction scenarios, account entities may encounter the following issues: 1) different accounts exhibit varying levels of interaction activity, leading to imbalanced degree distributions, as illustrated in Fig. 1; 2) the same account participates in different types of interactions to varying extents, resulting in imbalanced interaction distributions, as illustrated in Fig. 3. These diverse and imbalanced interaction phenomena can introduce biases when characterizing account interaction behaviors. To alleviate these issues, we introduce a generative self-supervised mechanism to augment the interaction features of Ethereum accounts, particularly for those with low activity levels and low-frequency interaction types.
@@ IV-A2 Feature Generation
where $\ast=\mathcal{T}_{\mathcal{V}}(v_{\textit{i}})$ indicates the type of $v_{\textit{i}}$, and $\hat{\boldsymbol{x}}_{\textit{i}}^{\textit{r}^{1}_{\ast}}$ is the feature generated conditional on $(\boldsymbol{x}_{\textit{i}},\boldsymbol{r}^{1}_{\ast})$. Note that under specific conditions, we can repeatedly sample the latent variable $\boldsymbol{z}$ to generate interaction features multiple times, i.e., repeatedly utilizing ICVAE to generate multiple interaction features under the same conditions, forming multiple feature views, and further alleviating the issue of feature distribution imbalance. The number of multiple feature views is denoted by $M$.
@@ IV-B Multi-view Ethereum Interaction Feature Learning
With the pre-trained ICVAE, we obtain augmented interaction features from multiple views, which can alleviate the issue of sparse interaction features for several accounts to some extent. To better characterize the account behavior patterns, the multi-view interaction feature learning module will process both the initial and generated interaction features to uncover potential fraud risks. This module will encode and fuse the multi-view interaction features and ultimately propagate the features by combining the structural information of HEIG. The design details of this module are described below.
@@ IV-B1 Input Feature Encoding
The multi-view interaction features generated by the ICVAE remain in the input space, with their dimensions represented as $\hat{\mathcal{X}}_{\textit{i}}\in\mathbb{R}^{M\times 4\times d}$. To further explore the underlying information in the interaction features, a common practice is to design a weight-sharing encoder for all target accounts, which accepts the interaction features of the accounts as input and outputs high-dimensional hidden features. However, we consider such practice to be inappropriate for the following reasons. We sample a certain number of EOA and CA and count their 14-dimensional features, as shown in Fig. 6, from which we can observe noticeable differences among the features of accounts within the same type, as well as among different types of accounts. In such cases, a weight-sharing transformation usually fails to adequately distinguish the differences among accounts. Therefore, to more adaptively characterize the interaction preferences of different types of accounts, we design account type-specific input feature encoding components, which follows a weight separation setting [40, 41]. Specifically, for the multi-view interaction features of target account $v_{\textit{i}}$ in the input space, we map them into the hidden space using an MLP parameterized by $\hat{\mathbf{\Theta}}_{\ast}\in\mathbb{R}^{d\times d^{\prime}}$, obtaining the hidden representation ^$\hat{\mathcal{H}}_{\textit{i}}\in\mathbb{R}^{M\times 4\times d^{\prime}}$ as follows:
@@ V-A Data Collection
We collect a certain amount of labeled data from the Xblock [^3] http://xblock.pro and Etherscan [^4] https://cn.etherscan.com platforms, covering 191 Ponzi accounts and 1151 non-Ponzi accounts, as well as 1206 phishing accounts and 1557 non-phishing accounts. Based on these labeled accounts, we further extract their first- and second-order transaction objects to construct the HEIG, obtaining two fraud detection datasets with imbalanced label distribution. The detailed statistics of these datasets is presence in Table I.
@@ V-B Comparison Methods
In this paper, we choose three categories of methods for comparison: machine learning-based method (LightGBM [21]), homogeneous graph analysis-based methods (Trans2Vec [22], GCN [26], GAT [29], SAGE [24], FA-GNN [28]), and heterogeneous graph analysis-based methods (HGT [42], RGCN [43], ieHGCN [44], HTSGCN [45]). We exclude metapath-based methods from our comparison due to the significant effort required in designing metapaths for targeting different fraud behaviors.
@@ V-C Experimental Settings
For all datasets, we divide them into training, validation, testing sets in a proportion of 6:2:2. All experiments will be run 5 times, with the average performance and standard deviation reported. Performance metrics used are Micro-F1 and Macro-F1. The former emphasizes overall performance across all categories, potentially obscuring the performance on minor categories. Conversely, the latter assigns equal weight to each category, thereby providing a measure of the ability to classify imbalanced categories. Simultaneously considering both metrics enables a comprehensive evaluation of detection performance across various categories.
TABLE II: The results of fraud detection in terms of Micro-F1(%), Macro-F1(%) and Standard Deviation(%). The best results are highlighted in bold, while the second-best results are underlined. “Rank” indicates the performance ranking of a method on a single fraud dataset, “Average Rank” indicates the average performance ranking of a method across all datasets, and “Group Rank” indicates the average performance ranking of a category of methods across all datasets.
Micro-F1 (%) Macro-F1 (%) Rank Micro-F1 (%) Macro-F1 (%) Rank
LightGBM 88.92$\pm$1.32 71.68$\pm$2.58 8.00 96.17$\pm$0.40 96.11$\pm$0.41 2.50 5.25 5.25
[…]
GCN 87.24$\pm$0.18 68.17$\pm$1.46 11.0 92.22$\pm$0.82 88.77$\pm$0.85 8.00 9.50
GAT 90.58$\pm$0.63 77.52$\pm$1.74 3.00 93.63$\pm$0.50 90.34$\pm$0.37 7.00 5.00
SAGE 89.71$\pm$0.35 73.91$\pm$1.11 5.50 95.01$\pm$0.62 91.69$\pm$0.60 5.50 5.50
FA-GNN 88.70$\pm$0.79 74.58$\pm$2.11 6.50 94.94$\pm$0.28 94.87$\pm$0.28 5.00 5.75
[…]
RGCN 89.67$\pm$1.01 75.79$\pm$2.39 5.00 95.29$\pm$0.57 92.06$\pm$0.45 4.50 4.75
ieHGCN 88.48$\pm$0.47 71.74$\pm$1.42 8.50 95.83$\pm$0.63 72.46$\pm$0.32 7.00 7.75
HTSGCN 88.17$\pm$0.28 71.35$\pm$1.31 10.0 96.49$\pm$0.56 72.93$\pm$0.12 5.50 7.75
[…]

### v2

@@ Abstract
The rampant fraudulent activities on Ethereum hinder the healthy development of the blockchain ecosystem, necessitating the reinforcement of regulations. However, multiple imbalances involving account interaction frequencies and interaction types in the Ethereum transaction environment pose significant challenges to data mining-based fraud detection research. To address this, we first propose the concept of meta-interactions to refine interaction behaviors in Ethereum, and based on this, we present a dual self-supervision enhanced Ethereum fraud detection framework, named Meta-IFD. This framework initially introduces a generative self-supervision mechanism to augment the interaction features of accounts, followed by a contrastive self-supervision mechanism to differentiate various behavior patterns, and ultimately characterizes the behavioral representations of accounts and mines potential fraud risks through multi-view interaction feature learning. Extensive experiments on real Ethereum datasets demonstrate the effectiveness and superiority of our framework in detecting common Ethereum fraud behaviors such as Ponzi schemes and phishing scams. Additionally, the generative module can effectively alleviate the interaction distribution imbalance in Ethereum data, while the contrastive module significantly enhances the framework’s ability to distinguish different behavior patterns. The source code will be available in https://github.com/GISec-Team/Meta-IFD.
@@ I Introduction
Figure 1: Illustration of the imbalanced distribution in account interaction frequency. Various types of accounts display long-tailed distributions for both types of fraud.
@@ II-A2 Phish Scam Detection
TABLE I: Notations and descriptions.
Notation Description
$G$ Heterogeneous Ethereum interaction graph
$\mathcal{V},\mathcal{E}$ Set of account nodes / interaction edges
$\mathcal{T}_{\mathcal{V}}$, $\mathcal{T}_{\mathcal{E}}$ Account / Interaction type mapping function
$\mathcal{R}$ Set of meta-interactions
$M$ Number of the views
$\boldsymbol{X}$ Account feature matrix
$\hat{\mathcal{X}}$ Set of augmented multi-view interaction features
$\hat{\mathcal{H}}$ Set of type-specific multi-view interaction features
$\bar{\mathcal{H}}$ Set of aggregated multi-view account features
$\bar{\boldsymbol{H}}$ Account feature matrix after multi-view fusing
$\boldsymbol{H}$ Final account representations
@@ II-B Graph-based Fraud Detection
Furthermore, the homogeneous transaction graphs constructed by the aforementioned methods neglect the heterogeneity of accounts and transaction behaviors. To address this, Zhou et al. [5] incorporate call records from contract accounts as features for externally owned accounts and propose a hierarchical attention encoder based on contrastive learning to effectively learn account representations, alleviating the issue of label scarcity. Du et al. [34] utilize a similar transaction graph construction method and design a GNN-based link prediction method to de-anonymize accounts in mixing services. Jin et al. [35, 36] predefine static and temporal meta-paths to describe the complex interaction behaviors in Ethereum and design a meta-path-based account feature augmentation module, which effectively improves the performance of existing Ponzi scheme detection methods. Considering that defining meta-paths requires expert knowledge and is time-consuming, Liu et al. [37] and Xu et al. [38] employ transformer models to autonomously acquire meta-path patterns, subsequently aggregating neighbor accounts based on these meta-paths to obtain higher-order information for enhancing fraud detection.
@@ III-C Refinement of Ethereum Interactions
Unlike the coarse-grained interactions trans and call which only reflect the ways in which accounts interact with each other, meta-interactions further describe the ways in which specific types of accounts can interact with each other, and thus can be referred to as fine-grained interactions. Based on the definition of meta-interaction and the characteristics of Ethereum, we can identify six types of meta-interactions from HEIG, as shown in Fig. 2(b). Furthermore, we define =$\mathcal{R}=\{r^{1}_{\text{ca}},r^{2}_{\text{ca}},r^{3}_{\text{ca}},r^{1}_{\text{eoa}},r^{2}_{\text{eoa}},r^{3}_{\text{eoa}}\}$ to denote the set of meta-interactions, and the detailed description is shown in Table. II.
TABLE II: Descriptions of different meta-interactions.
Meta-interactions Descriptions
$r^{1}_{\text{ca}}:\text{CA}\stackrel{{\scriptstyle\text{{call}}}}{{\longrightarrow}}\text{CA}$ Interaction between CA to achieve complex logical functions or extend functionalities.
$r^{1}_{\text{eoa}}:\text{EOA}\stackrel{{\scriptstyle\text{{call}}}}{{\longrightarrow}}\text{CA}$ EOA initiates a request to call contract functions or update contract states.
$r^{2}_{\text{ca}}:\text{CA}\stackrel{{\scriptstyle\text{{trans}}}}{{\longrightarrow}}\text{CA}$ CA handles automated asset management based on predefined internal logic.
$r^{2}_{\text{eoa}}:\text{EOA}\stackrel{{\scriptstyle\text{{trans}}}}{{\longrightarrow}}\text{CA}$ EOA transfers Ether or tokens to CA to provide funding or trigger its execution logic.
$r^{3}_{\text{ca}}:\text{CA}\stackrel{{\scriptstyle\text{{trans}}}}{{\longrightarrow}}\text{EOA}$ CA transfers Ether or tokens to an EOA, distributing funds or rewards.
$r^{3}_{\text{eoa}}:\text{EOA}\stackrel{{\scriptstyle\text{{trans}}}}{{\longrightarrow}}\text{EOA}$ Direct Ether or token transfers between EOAs without involving smart contract logic.
@@ IV-A Fine-grained Ethereum Interaction Feature Generation
In complex Ethereum interaction scenarios, account entities may encounter the following issues: 1) different accounts exhibit varying levels of interaction activity, leading to imbalanced degree distributions, as illustrated in Fig. 1, which reveals typical long-tail distributions in both datasets. Specifically, the number of low-degree (inactive) accounts significantly exceeds that of high-degree (active) accounts, thus forming the classical degree imbalanced problem in complex networks; 2) the same account participates in different types of interactions to varying extents, resulting in imbalanced interaction distributions, as illustrated in Fig. 4, where box plot represents the statistical distribution of account degrees across various interactions, including the minimum, upper quartile, median, lower quartile and the maximum value. Long and thin boxes indicate a large difference in the degree distribution of accounts under this interaction type. The variations in box shapes between different interaction types suggest greater differences in degree distributions between interaction types. These diverse and imbalanced interaction phenomena can introduce biases when characterizing account interaction behaviors. To alleviate these issues, we introduce a generative self-supervised mechanism to augment the interaction features of Ethereum accounts, particularly for those with low activity levels and low-frequency interaction types.
Figure 4: Illustration of the imbalanced distribution in account interaction types. Boxes indicate the degree correlation statistics for labeled accounts across different interaction types.
@@ IV-A1 Model Architecture and Pre-training
Algorithm 1 The training process of Meta-IFD.
0:  Heterogeneous Ethereum interaction graph $G=(\mathcal{V},\mathcal{E},\mathcal{T}_{\mathcal{V}},\mathcal{T}_{\mathcal{E}},\boldsymbol{X},\mathcal{Y})$, meta-interaction set $\mathcal{R}$, the number of views $M$, and pre-trained ICVAE.
0:  Meta-IFD.
1:  Load pre-trained ICVAE model;
2:  for each account $v_{\textit{i}}\in\mathcal{V}$ do
3:     Get feature vector $\boldsymbol{x}_{\textit{i}}$;
4:     while $M>0$ do
5:        for each neighbor account $v_{\textit{j}}\in\mathcal{N}_{v_{\textit{i}}}$ do
6:           for different meta-interaction $r_{\textit{ij}}\in\mathcal{R}$ do
7:              Generate interaction features $\hat{\boldsymbol{x}}_{\textit{i}}$ via Eq. (4);
8:           end for
9:        end for
10:        $M\leftarrow M-1$;
11:     end while
12:  end for
13:  Obtain augmented interaction feature $\hat{\mathcal{X}}$;
14:  Perform type-specific transformation via Eq. (6);
15:  Calculate coarse-grained interaction feature contrast loss $\mathcal{L}_{\text{con}}^{\ast}$ via Eq. (11)
16:  Perform intra-view feature aggregation via Eq. (7);
17:  Perform inter-view feature aggregation via Eq. (9);
18:  Obtaining final account representation $\boldsymbol{H}$ with two-layer message passing according to Eq. (10);
19:  Calculate classification loss via Eq. (12);
20:  Train Meta-IFD via minimize the optimization objective in Eq. (13);
21:  return  Meta-IFD.
@@ IV-A2 Feature Generation
where $\ast=\mathcal{T}_{\mathcal{V}}(v_{\textit{i}})$ indicates the type of $v_{\textit{i}}$, and $\hat{\boldsymbol{x}}_{\textit{i}}^{\textit{r}^{1}_{\ast}}$ is the feature generated conditional on $(\boldsymbol{x}_{\textit{i}},\boldsymbol{r}^{1}_{\ast})$. Note that under specific conditions, we can repeatedly sample the latent variable $\boldsymbol{z}$ to generate interaction features multiple times, i.e., repeatedly utilizing ICVAE to generate multiple interaction features under the same conditions, forming multiple feature views, and further alleviating the feature distribution imbalance issue. The number of multiple feature views is denoted by $M$.
@@ IV-B Multi-view Ethereum Interaction Feature Learning
With the pre-trained ICVAE, we obtain augmented interaction features from multiple views, which can alleviate the issue of sparse interaction features for several accounts to some extent. To better characterize the account behavior patterns, the multi-view interaction feature learning module will process both the initial and generated interaction features to uncover potential fraud risks. This module will encode and fuse the multi-view interaction features and ultimately propagate the features by combining the structural information of HEIG. The design details of the module are described as follows, and the complete process is presented in Algorithm 1.
@@ IV-B1 Input Feature Encoding
The multi-view interaction features generated by the ICVAE remain in the input space, with their dimensions represented as $\hat{\mathcal{X}}_{\textit{i}}\in\mathbb{R}^{M\times 4\times d}$. To further explore the underlying information in the interaction features, a common practice is to design a weight-sharing encoder for all target accounts, which accepts the interaction features of the accounts as input and outputs high-dimensional hidden features. However, we consider such practice to be inappropriate for the following reasons. We sample a certain number of EOA and CA and count their 14-dimensional normalized features, as shown in Fig. 6, where the first 7-dimensional features are derived from the call, and the last 7-dimensional features are derived from the trans. We can observe noticeable differences among the features of accounts within the same type, as well as among different types of accounts. For example, there is a significant difference in the call features extracted from EOA and CA, and the trans features of EOA are relatively larger than those of CA. In such cases, a weight-sharing transformation usually fails to adequately distinguish the differences among accounts. Therefore, to more adaptively characterize the interaction preferences of different types of accounts, we design account type-specific input feature encoding components, which follows a weight separation setting [40, 41]. Specifically, for the multi-view interaction features of target account $v_{\textit{i}}$ in the input space, we map them into the hidden space using an MLP parameterized by $\hat{\mathbf{\Theta}}_{\ast}\in\mathbb{R}^{d\times d^{\prime}}$, obtaining the hidden representation ^$\hat{\mathcal{H}}_{\textit{i}}\in\mathbb{R}^{M\times 4\times d^{\prime}}$ as follows:
@@ V-A Data Collection
We collect a certain amount of labeled data from the Xblock [^3] http://xblock.pro and Etherscan [^4] https://cn.etherscan.com platforms, covering 191 Ponzi accounts and 1151 non-Ponzi accounts, as well as 1206 phishing accounts and 1557 non-phishing accounts. Based on these labeled accounts, we further extract their first- and second-order transaction objects to construct the HEIG, obtaining two fraud detection datasets with imbalanced label distribution. The detailed statistics of these datasets is presence in Table III. According to the statistics, we observe that although the Ponzi dataset contains fewer labeled instances, it exhibits significantly richer interaction behaviors. Specifically, the total number of interactions in the Ponzi dataset exceeds that of the Phish dataset by 15,229,984, with an additional 8,192,525 call interactions. This discrepancy arises because executing Ponzi fraud requires multiple contract triggers, whereas phishing fraud in the Phish dataset is typically completed by merely guiding users to perform a single transfer.
@@ V-B Comparison Methods
In this paper, we choose four categories of methods for comparison: machine learning-based method (LightGBM [21]), homogeneous graph analysis-based methods (Trans2Vec [22], GCN [26], GAT [29], SAGE [24], FA-GNN [28]), heterogeneous graph analysis-based methods (HGT [42], RGCN [43], ieHGCN [44], HTSGCN [45]), and sequence analysis-based methods (BERT4ETH [18], ZipZap [19]). We exclude meta-path-based methods from our comparison due to the significant effort required in designing meta-paths for targeting different fraud behaviors.
TABLE IV: The results of fraud detection in terms of Precision(%), Recall(%), Binary-F1(%), Micro-F1(%), Macro-F1(%) and Standard Deviation(%). The best results are highlighted in bold, while the second-best results are underlined. “Rank” indicates the performance ranking of a method on a single fraud dataset, “Average Rank” indicates the average performance ranking across all datasets, and “Group Rank” indicates the average performance ranking of a category of methods across all datasets.
Precision Recall Binary-F1 Micro-F1 Macro-F1 Rank Precision Recall Binary-F1 Micro-F1 Macro-F1 Rank
LightGBM 72.29$\pm$11.91 38.42$\pm$2.11 49.97$\pm$4.18 89.07 $\pm$1.30 71.92$\pm$2.46 8.2 94.13$\pm$ 0.67 97.01$\pm$1.13 95.54$\pm$0.37 96.06 $\pm$0.31 96.00$\pm$0.32 4.8 6.5 6.50
[…]
GCN 67.61$\pm$8.70 27.89$\pm$6.78 38.44$\pm$5.22 87.66$\pm$0.49 65.78$\pm$2.51 12.2 92.05$\pm$1.35 88.38$\pm$1.14 90.18$\pm$1.19 91.61$\pm$1.03 91.43$\pm$1.05 11.6 11.9
GAT 77.62$\pm$6.48 47.37$\pm$6.66 58.40$\pm$5.01 90.56$\pm$0.80 76.53$\pm$2.68 5.0 90.20$\pm$1.52 95.19$\pm$0.86 92.62$\pm$0.95 93.38$\pm$0.90 93.31$\pm$0.90 9.6 7.3
SAGE 70.88$\pm$7.06 37.37$\pm$3.07 48.78$\pm$3.26 88.92$\pm$0.79 71.29$\pm$1.82 9.6 93.46$\pm$0.45 96.02$\pm$0.62 94.72$\pm$0.44 95.33$\pm$0.39 95.27$\pm$0.39 6.2 7.9
FA-GNN 66.69$\pm$6.00 42.11$\pm$4.40 51.17$\pm$1.76 88.70$\pm$0.50 72.39$\pm$0.76 9.0 96.92$\pm$0.20 96.60$\pm$0.31 96.76$\pm$0.21 97.18$\pm$0.18 97.13$\pm$0.19 2.6 5.8
[…]
RGCN 69.61$\pm$5.64 53.16$\pm$5.10 60.26$\pm$5.29 90.11$\pm$1.26 77.31$\pm$3.00 5.6 94.95$\pm$1.57 94.36$\pm$1.07 94.64$\pm$0.45 95.33$\pm$0.43 95.25$\pm$0.43 6.4 6.0
ieHGCN 61.49$\pm$5.34 38.42$\pm$11.36 46.59$\pm$10.69 88.03$\pm$1.21 69.92$\pm$5.62 11.4 94.38$\pm$0.83 94.36$\pm$3.63 94.32$\pm$1.75 95.08$\pm$1.39 94.99$\pm$1.45 7.4 9.4
HTSGCN 67.47$\pm$5.79 42.11$\pm$3.33 51.59$\pm$2.34 88.85$\pm$0.78 72.64$\pm$1.30 8.2 97.00$\pm$0.68 96.27$\pm$0.95 96.63$\pm$0.37 97.07$\pm$0.31 97.02$\pm$0.32 3.2 5.8
[…]
ZipZap 69.64$\pm$8.44 56.22$\pm$4.65 61.77$\pm$3.68 90.37$\pm$1.28 78.13$\pm$2.17 4.6 98.27$\pm$1.24 94.22$\pm$1.61 96.19$\pm$0.47 96.96$\pm$0.35 96.83$\pm$0.37 4.4 4.5
[…]
@@ V-C Experimental Settings
For sequence analysis-based methods, during the pre-trained phase, we utilize default parameters with Transformer layers set to 8, hidden dimension set to 64, number of attention heads set to 2, maximum sequence length set to 100, and the optimal learning rate selected from {0.001, 0.0001}. During the classification phase, a two-layer MLP is employed with an optimal combination of learning rate chosen from {0.01, 0.001} and hidden dimensions selected from {32, 64, 128}, respectively. The parameter configuration ensures that the search ranges for hyperparameters align with those employed in the comparative detection methods discussed in this paper, thereby maintaining the fairness of the experimental comparison.
For all datasets, we divide them into training, validation, testing sets in a proportion of 6:2:2. To prevent overfitting, we set an early stop with a patience of 50. All experiments will be run 5 times, with the average performance and standard deviation reported. Performance metrics used are Precision, Recall, Binary-F1, Micro-F1 and Macro-F1. The first three metrics evaluate the detection performance for positive samples with the binary evaluation. Micro-F1 emphasizes the overall performance across all categories, akin to Accuracy in binary classification. Macro-F1 calculates the Binary-F1 metrics of both positive and negative samples separately and reports the average results. Combining the above metrics, we can evaluate the model’s detection ability for each sample type.